A 25.0 mL solution of HNO3 is neutralized with 22.4 mL of 0.250 M Ba (OH)2. What is the concentration of the original HNO3 solution?
1
Expert's answer
2020-04-28T12:21:35-0400

"Ba(OH)_2+2HNO_3\to Ba(NO_3)_2+2H_2O"

As per the reaction,

"2\times" Moles of "Ba(OH)_2" reacted"=" Moles of "HNO_3" reacted.

And,Molarity equation for the reaction will be:

"\frac{M_1V_1}{n_1}=\frac{M_2V_2}{n_2}" ,where "M,V,n" represent concentration,volume and stochiometric cofficients for each of the reactants and "1" represent for "Ba(OH)_2" and "2" for "HNO_3."

"\implies \frac{22.4\times0.250}{1}=\frac{25\times\ M_2}{2}"

"\implies M_2=0.448\ M=" Concentration of the original "HNO_3" solution